# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ Abstract
1-identification is a fundamental multi-armed bandit formulation on pure exploration. An agent aims to determine whether there exists a qualified arm whose mean reward is not less than a known threshold $\mu_{0}$, or to output None if it believes such an arm does not exist. The agent needs to guarantee its output is correct with probability at least $1-\delta$, while making expected total pulling times $\mathbb{E}\tau$ as small as possible. We work on 1-identification with two main contributions. (1) We utilize an optimization formulation to derive a new lower bound of $\mathbb{E}\tau$, when there is at least one qualified arm. (2) We design a new algorithm, deriving tight upper bounds whose gap to lower bounds are up to a polynomial of logarithm factor across all problem instance. Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem left by history literature.
Keywords 1-identification, Fixed Confidence, Pure Exploration
@@ 1.1 Model & Problem Formulation
We study the 1-identification problem [^1] This name origins from Katz-Samuels and Jamieson (2020), Li and Cheung (2025). In Degenne and Koolen (2019), it is also called Any Low Problem., a variant of the best arm identification problem. An instance of 1-identification consists of a tuple of four elements: $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$. The set $[K]$ denotes the arm set $\{1,2,\cdots,K\}$, and we assume $K\geq 2$. For each $a\in[K]$, $\rho_{a}$ is the reward probability distribution associated with arm $a$. The reward distribution $\rho_{a}$ is not known to the agent for each $a$. By pulling arm $a$ once, the agent receives a random reward distributed as $\rho_{a}$. We denote $r_{a}^{\nu}:=\mathbb{E}_{R\sim\rho_{a}}[R]$ as the mean reward of arm $a$ under reward distribution $\rho_{a}$. We assume the random noise $R_{a}-r_{a}^{\nu}$ to be 1-sub-Gaussian. The real number $\mu_{0}$ is a threshold known to the agent. The agent’s main goal is to predict whether there is an arm $a\in[K]$ with $r_{a}^{\nu}\geq\mu_{0}$, and to identify such an arm in the positive case, while outputting “None” in the negative case. The quantity $\delta\in(0,1)$ is the tolerant parameter on the probability of a wrong prediction.
The agent’s pulling strategy is specified by a sampling rule $\pi=\{\pi_{t}\}_{t=1}^{\infty}$, a stopping time $\tau$ and a recommendation rule $\hat{a}\in[K]\cup\{\textsf{None}\}$. When a pulling strategy is applied on a 1-identification problem instance with reward distribution $\rho$, the strategy generates a history $\{U_{0}\}\cup\{A_{t},X_{t},U_{t}\}_{t=1}^{\tau}$, where $A_{t}\in[K]$ denotes the arm pulled at $t$, $X_{t}\sim\rho_{A_{t}}$ denotes the reward received, and $U_{t}\sim\text{Unif}(0,1)$ is an internal random seed. Arm $A_{t}\in[K]$ is chosen contingent upon the history $H(t-1)=\{U_{0}\}\cup\{(A_{s},X_{s},U_{s})\}^{t-1}_{s=1}$, via the function $\pi_{t}(H(t-1))$. The agent stops pulling at the end of time step $\tau$, where $\tau$ is a stopping time with respect to the filtration $\{\sigma(H(t))\}^{\infty}_{t=1}$. After time step $\tau$, the agent outputs the prediction $\hat{a}\in[K]\cup\{\textsf{None}\}$. We require that $\hat{a}$ is measurable with respect to $\sigma(H(\tau))$, i.e. determined by the information $H(\tau)$. Outputting $\hat{a}\in[K]$ means that the agent concludes with arm $\hat{a}$ satisfying $r_{\hat{a}}^{\nu}\geq\mu_{0}$, while outputting $\hat{a}=\textsf{None}$ means that the agent concludes with no arm having mean reward $\geq\mu_{0}$. We say that $\nu$ is a positive instance if $\max_{a}r_{a}^{\nu}>\mu_{0}$, and say that $\nu$ is a negative instance if $\max_{a}r_{a}^{\nu}<\mu_{0}$.
For $\Delta>0$, we define $\mathcal{S}^{\text{pos}}_{\Delta}=\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\}$, $\mathcal{S}^{\text{neg}}_{\Delta}=\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\}$, and take $\mathcal{S}^{\text{pos}}=\cup_{\Delta>0}\mathcal{S}^{\text{pos}}_{\Delta},\mathcal{S}^{\text{neg}}=\cup_{\Delta>0}\mathcal{S}^{\text{neg}}_{\Delta}$. We define $i^{*}(\nu)$ as the set of correct answers. For $\nu\in\mathcal{S}^{\text{pos}}_{\Delta}$, we have $i^{*}(\nu)=\{a:r_{a}^{\nu}\geq\mu_{0}\}$, while for instance $\nu\in\mathcal{S}^{\text{neg}}_{\Delta}$, we have $i^{*}(\nu)=\{\textsf{None}\}$. We don’t consider instance whose $\max r_{a}^{\nu}=\mu_{0}$ and mainly focus on analyzing PAC algorithms, with the following definitions.
Definition 1.
A pulling strategy is $\delta$-PAC, if for any $\delta\in(0,1),\nu\in\mathcal{S}^{\text{pos}}\cup\mathcal{S}^{\text{neg}}$, it satisfies $\Pr_{\nu}(\tau<+\infty,\hat{a}\in i^{*}(\nu))>1-\delta$.
Appendix A.3 in Li and Cheung (2025) shows that a $\delta$-PAC algorithm can still suffer from $\mathbb{E}_{\nu}\tau=+\infty$ for some $\nu\in\mathcal{S}^{\text{pos}}$. Thus, we further introduce the following definition, which can be considered “Stronger $\delta$-PAC”.
Definition 2.
A pulling strategy is $(\Delta,\delta)$-PAC, if it is $\delta$-PAC, and for any $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\text{pos}}_{\Delta}\cup\mathcal{S}^{\text{neg}}_{\Delta}}\mathbb{E}_{\nu}\tau<+\infty$.
@@
1.2 Lit Review on 1-identification
Before illustrating our main contribution, we first review the weakness in the history literature and we start the review by unifying the notations. In the following paper, we consider a reward vector $\{\mu_{a}\}_{a=1}^{K}\subset[0,1]^{K}$ satisfies that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$. Also, we consider $\mu_{0}\in[0,1]$ and define $\Delta_{i,j}=|\mu_{i}-\mu_{j}|$ for $\forall i,j\in[K]\cup\{0\}$. For an instance $\nu$, we assume $r_{a}^{\nu}=\mu_{a}$ without further explanation. In addition, we introduce different complexity terms occurred in the history result.
@@ 1.2 Lit Review on 1-identification
The discussion on the sample complexity focuses on two cases: instance $\nu$ is negative or positive. There has been abundant research achieving the tight approximation of sample complexity on a negative instance, which is similar to the Best Arm Identification. Regarding the lower bound of $\mathbb{E}\tau$, Theorem 5.4 in Li and Cheung (2025) asserts $\mathbb{E}\tau\geq\Omega(H_{1}^{\text{neg}}\log\frac{1}{\delta})$, if we apply any $\delta$-PAC algorithm on an instance $\nu$. Meanwhile, Algorithm HDoC in Kano et al. (2017), APGAI in Jourdan and Réda (2023), SEE in Li and Cheung (2025), are able to guarantee $\mathbb{E}\tau\leq O(H_{1}^{\text{neg}}(\log\frac{1}{\delta}+\log H_{1}^{\text{neg}}))$, ignoring some logarithm factor. In addition, Algorithm Sticky Track and Stop (S-TaS) in Degenne and Koolen (2019) is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\text{neg}}$ if $\nu$ is negative and the reward follows Gaussian distribution with variance $1$.
Contrast with the negative instance, the complexity of positive instance remains a mystery, as the proposed upper bounds for a positive instance suffer from different looseness. For simplicity, we consider a positive instance $\nu$ and denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and only consider reward following Gaussian distribution with variance $1$. We propose Table 1 to summarize the looseness in the current result, with imprecise formulation because of space limit. The actual upper bound existing in the history literature is not smaller than our presented formulation, and we can still observe the looseness in our presented formulation. Following the Appendix A.2 in Li and Cheung (2025), we put details of simplification on history upper bound in Appendix A.
@@
S-TaS, Degenne and Koolen (2019)
Poiani et al. (2025)
@@
Parallel-SEE, This work $\mathbb{E}\tau\leq O\Big(\log^{4}K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log\frac{1}{\Delta_{j,0}^{2}}\Big)$ $\surd$
@@
LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$ $\surd$
Degenne and Koolen (2019) propose algorithm S-TaS, which provides asymptotic guarantee for both negative and positive instances. S-TaS is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq\frac{2}{\Delta_{0,1}^{2}}$. Together with their proposed lower bound $\lim\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\geq\frac{2}{\Delta_{0,1}^{2}}$, we know S-TaS is asymptotically optimal. Degenne and Koolen (2019) leaves non-asymptotic analysis of S-TaS as a future direction, which is complemented by Poiani et al. (2025). Poiani et al. (2025) present the non-asymptotic upper bound in an implicit way, but we are able to derive a smaller but explicit formulation. The weakness of Poiani et al. (2025) comes from the $\delta$-independent part $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which is inconsistent with any lower bound.
@@
Algorithm HDoC in Kano et al. (2017) proves an upper bound not smaller than $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$. The suboptimality of $C^{\text{HDoC}}_{\text{pos}}$ consists of two parts. The first part is $H_{1}\log\log\frac{1}{\delta}$, with comments in Kano et al. (2017): “Therefore, in the case that $\ldots$ $K=\Omega(\frac{\log\frac{1}{\delta}}{\log\log\frac{1}{\delta}})$there still exists a gap between the lower bound in …and the upper bound in …”. The second part is $\frac{K}{\epsilon^{2}}$, which can be infinity if $\mu_{i}=\mu_{j}$ holds. for some $i,j\in[K]$. Based on HDoC, Tsai et al. (2024) propose Algorithm lilHDoC, which incorporate the LIL concentration event (Lemma 3 in Jamieson et al. (2014)) and add an extra warm-up stage. After adding these designs, the proposed algorithm lilHDoC still suffers from the same sub-optimality as HDoC.
Kano et al. (2017) design Algorithm APGAI to address Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$. Though it is not the main focus, they also propose an upper bound for 1-identification. The weakness of upper bound $O(H_{0}(\log\frac{K}{\delta}))$ mainly comes from $H_{0}$. Easy to see $H_{0}$ is proportional to $K$, suggesting that the coefficient of $\log\frac{1}{\delta}$ is suboptimal compared to the others. What’s more, if there exists an arm $a$ such that $\mu_{a}=\mu_{0}$, we have $H_{0}=\infty$, suggesting that the upper bound is vacuous.
Katz-Samuels and Jamieson (2020) work on $\epsilon$-good identification and 1-identification simultaneously. They define a new concept Unverifiable Sample Complexity and prove upper bounds for it. Though Unverifiable Sample Complexity is different with Sample Complexity in this paper, we can still adapt their result, as discussed by Appendix A.1 in Li and Cheung (2025). The suboptimality of its upper bound consists of two points. First, its upper bound relies on $\log\frac{1}{\delta}\log\log\frac{1}{\delta}$ instead of $\log\frac{1}{\delta}$. Second, the complexity term $H^{u}$ relies on the minimum gap between a qualified mean reward and the threshold $\mu_{0}$, which can be much smaller than $\Delta_{0,1}$. In the case $m$ is much smaller than $K$ and $\Delta<\Delta_{0,1}$, this upper bound is much larger than their proposed lower bound.
@@
Li and Cheung (2025) design Algorithm SEE, splitting the upper bound of $\mathbb{E}\tau$ into two parts which are $\delta$-dependent and $\delta$-independent. They prove both upper and lower bounds of $\mathbb{E}\tau$ with explicit formulation. They show when $m=1$, i.e. there is only one arm whose reward is greater than $\mu_{0}$, the gap between their proposed upper and lower bounds are up to a polynomial of $\log K,\log\frac{1}{\Delta_{0,1}^{2}}$, suggesting the optimality. However, they also comment $H(1)\log\frac{K}{\Delta_{0,1}^{2}}$, which is $\delta$-independent part, might be improved if $m>1$. They leave the possible improvement as an open direction.

### v2

@@ Abstract
The 1-identification problem is a fundamental pure-exploration problem in multi-armed bandits. An agent aims to determine whether there exists an arm whose mean reward exceeds a known threshold $\mu_{0}$, or to output None otherwise. The agent must guarantee correctness with probability at least $1-\delta$, while minimizing the expected number of arm pulls $\mathbb{E}[\tau]$. We study the 1-identification problem and make two main contributions. First, for instances with at least one qualified arm, we derive a new lower bound on $\mathbb{E}[\tau]$ via a novel optimization formulation. Second, we propose a new algorithm and establish upper bounds that match the lower bounds up to polynomial logarithmic factors uniformly over all instances. Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem in the literature.
@@ 1.1 Model & Problem Formulation
We study the 1-identification problem [^1] The terminology originates from Katz-Samuels and Jamieson (2020); Li and Cheung (2025). Degenne and Koolen (2019) call it Any Low Problem., a variant of the best arm identification problem. An instance of 1-identification is specified by the tuple $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$, where $K]\coloneqq\{1,2,\ldots,K\}$ denotes the set of arms with $K\geq 2$, and $\rho_{a}$ is the unknown reward distribution associated with arm $a$. Upon pulling arm $a$, the agent observes a random reward $R\sim\rho_{a}$. We denote by $r_{a}^{\nu}\coloneqq\mathbb{E}_{R\sim\rho_{a}}[R]$ the mean reward of arm $a$ under instance $\nu$, and assume that the noise $R-r_{a}^{\nu}$ is $1$-sub-Gaussian. The threshold $\mu_{0}\in\mathbb{R}$ is known to the agent. The confidence parameter $\delta\in(0,1)$ controls the allowable probability of an incorrect decision. The objective of the agent is to determine whether there exists an arm $a\in[K]$ such that $r_{a}^{\nu}\geq\mu_{0}$. In the affirmative case, the agent must identify (at least) one such arm; otherwise, it must output the symbol None.
The 1-identification problem serves as a framework for many decision-making scenarios. For instance, a clinical trial screening phase where researchers must decide whether any of several new drug candidates meet a specific efficacy threshold to justify the cost of further development. Similarly, considering anomaly detection in power grids, an agent may need to determine if if any substation is experiencing a voltage fluctuation that exceeds safety limits. In both cases, the agent does not necessarily need to rank every option; rather, it must quickly identify at least one "qualified" arm or confidently conclude that none exist.
@@
Instance classes. Following Li and Cheung (2025), an instance $\nu$ is called a positive instance if $\max_{a\in[K]}r_{a}^{\nu}>\mu_{0}$, and a negative instance if $\max_{a\in[K]}r_{a}^{\nu}<\mu_{0}$. For $\Delta>0$, we define the margin-separated classes $\mathcal{S}^{\mathrm{pos}}_{\Delta}:=\bigl\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\bigr\}$, $\mathcal{S}^{\mathrm{neg}}_{\Delta}\coloneqq\bigl\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\bigr\}$. Let $\mathcal{S}^{\mathrm{pos}}\coloneqq\bigcup_{\Delta>0}\mathcal{S}^{\mathrm{pos}}_{\Delta}$, $\mathcal{S}^{\mathrm{neg}}\coloneqq\bigcup_{\Delta>0}\mathcal{S}^{\mathrm{neg}}_{\Delta}$. We define the set of correct answers $i^{*}(\nu)$ by $i^{*}(\nu)\coloneqq\begin{cases}\{a\in[K]:r_{a}^{\nu}\geq\mu_{0}\},&\text{if }\nu\in\mathcal{S}^{\mathrm{pos}},\\[4.0pt] \{\textsf{None}\},&\text{if }\nu\in\mathcal{S}^{\mathrm{neg}}.\end{cases}$. In this work, We ignore instances for which $\max_{a}r_{a}^{\nu}=\mu_{0}$.
PAC guarantees. We adopt a PAC framework and focus on algorithms with correctness guarantees.
Definition 1 ($\delta$-PAC).
An algorithm is $\delta$-PAC, if for any $\nu\in\mathcal{S}^{\mathrm{pos}}\cup\mathcal{S}^{\mathrm{neg}}$, $\Pr_{\nu}\!\left(\tau<\infty,\;\hat{a}\in i^{*}(\nu)\right)\geq 1-\delta.$
As shown in Appendix A.3 in Li and Cheung (2025), a $\delta$-PAC algorithm may still satisfy $\mathbb{E}_{\nu}[\tau]=\infty$ for some $\nu\in\mathcal{S}^{\mathrm{pos}}$. To rule out such pathological behavior, we introduce a stronger notion.
Definition 2 (Strong $\delta$-PAC).
An algorithm is Strong $\delta$-PAC if it is $\delta$-PAC and, for all $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\mathrm{pos}}_{\Delta}\cup\mathcal{S}^{\mathrm{neg}}_{\Delta}}\mathbb{E}_{\nu}[\tau]<\infty$.
[…]
1.2 Literature Review
Before illustrating our main contribution, we review the existing literature on the 1-identification problem and highlight its current limitations. We begin by unifying notation across prior works in order to facilitate comparison. Throughout this paper, we consider a reward vector $(\mu_{a})_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$, together with a known threshold $\mu_{0}\in[0,1]$. For any $i,j\in[K]\cup\{0\}$, we define the gap $\Delta_{i,j}\coloneqq|\mu_{i}-\mu_{j}|$. In the analysis of upper bound(section 3), we only consider instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$. [^2] The agent does not know the monotonic ordering of $\{\mu_{a}\}_{a=1}^{K}$, and this default setting is for convenience in discussions. In the analysis of lower bound(section 2), we further consider instances $\nu^{\prime}$ whose reward vector $r_{a}^{\nu^{\prime}}$ is not ordered. We next recall several instance-dependent complexity measures that appear in the literature:
@@ 1.2 Literature Review
The study of sampling complexity naturally separates into two regimes, depending on the instance $\nu$ is negative or positive. For negative instances, the problem is closely related to the classical Best-Arm Identification problem, and the sample complexity is well studied. In particular, Theorem 5.4 of Li and Cheung (2025) establishes that, for any $\delta$-PAC algorithm and any negative instance $\nu$, $\mathbb{E}_{\nu}[\tau]\geq\Omega\!\left(H_{1}^{\mathrm{neg}}\log\frac{1}{\delta}\right)$. On the algorithmic side, HDoC Kano et al. (2017), APGAI Jourdan and Réda (2023), and SEE Li and Cheung (2025) guarantee $\mathbb{E}_{\nu}[\tau]\leq O(H_{1}^{\mathrm{neg}}(\log(1/\delta)+\log H_{1}^{\mathrm{neg}}))$, up to logarithmic factors. Moreover, for Gaussian rewards with unit variance, the Sticky Track-and-Stop (S-TaS) algorithm Degenne and Koolen (2019) achieves the asymptotically optimal guarantee $\limsup_{\delta\to 0}\frac{\mathbb{E}_{\nu}[\tau]}{\log(1/\delta)}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\mathrm{neg}}$ for negative instances.
In contrast, the sample complexity of positive instances remains unclear. Existing upper bounds suffer from various looseness and fail to match known lower bounds except in restricted instances. To illustrate with simplicity, we consider a positive instance $\nu$ satisfying $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and restrict attention to Gaussian rewards with variance 1. Table 1 summarizes representative non-asymptotic upper and lower bounds from the literature. Due to space limit, the bounds in the table are simplified; the original bounds are larger than those reported here. Even under these simplifications, the looseness remains apparent. Following Appendix A.2 of Li and Cheung (2025), we provide a detailed derivation of the simplified bounds in Appendix A.
@@
S-TaS,
proposed by Degenne and Koolen (2019)
analyzed by Poiani et al. (2025)
@@
PSEEB, This work $\mathbb{E}\tau\leq O\Big(\log K\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log^{4}K\log\frac{1}{\Delta_{j,0}^{2}}\right\}\Big)$ $\surd$
@@
LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$ $\surd$
Degenne and Koolen (2019) propose the S-TaS algorithm, which attains asymptotically optimal guarantees for both negative and positive instances. In particular, they show $\limsup_{\delta\to 0}\frac{\mathbb{E}_{\nu}[\tau]}{\log(1/\delta)}\leq\frac{2}{\Delta_{0,1}^{2}}$, together with a matching lower bound $\lim\inf_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log(1/\delta)}\geq\frac{2}{\Delta_{0,1}^{2}}$, establishing asymptotic optimality. A non-asymptotic analysis is left as a future direction in Degenne and Koolen (2019) and later addressed by Poiani et al. (2025). The resulting non-asymptotic bound, however, contains a $\delta$-independent term $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which has no counterpart in known lower bounds and leads to looseness.
@@
The HDoC algorithm of Kano et al. (2017) yields an upper bound $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\{\min_{a}\Delta_{0,a},\,\min_{a}\Delta_{a,a+1}/2\}$. Its suboptimality arises from two sources. The first source is $H_{1}\log\log\frac{1}{\delta}$, which is explicitly acknowledged in Kano et al. (2017): “Therefore, in the case that $\ldots$ $K=\Omega(\frac{\log\frac{1}{\delta}}{\log\log\frac{1}{\delta}})$there still exists a gap between the lower bound in …and the upper bound in …”. The second source is $\frac{K}{\epsilon^{2}}$, which is infinite when $\epsilon=0$. The lilHDoC algorithm Tsai et al. (2024), despite incorporating LIL-based confidence bounds and an additional warm-up phase, inherits the same fundamental limitations.
The APGAI algorithm Jourdan and Réda (2023) is originally designed for Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$. But the analysis in Jourdan and Réda (2023) also implies an upper bound $O(H_{0}\log(K/\delta))$ for 1-identification. This bound is loose primarily due to the dependence on $H_{0}$, which might scale linearly with $K$, leading to a suboptimal coefficient of $\log(1/\delta)$. Moreover, if there exists an arm with $\mu_{a}=\mu_{0}$, then $H_{0}=\infty$. In this case, the bound is vacuous.
Katz-Samuels and Jamieson (2020) study $\epsilon$-good identification and 1-identification via the notion of Unverifiable Sample Complexity. Although this notion differs from the sampling complexity considered here, their results can be adapted as discussed in Appendix A.1 of Li and Cheung (2025). The resulting upper bound is suboptimal for two reasons: it scales as $\log(1/\delta)\log\log(1/\delta)$ rather than $\log(1/\delta)$, and it depends on the complexity term $H^{\mathrm{u}}$, which is governed by the smallest gap above $\mu_{0}$ and may be much smaller than $\Delta_{0,1}$. When $m\ll K$ and $\Delta\ll\Delta_{0,1}$, this leads to a significant gap compared to the lower bound.
@@
Li and Cheung (2025) propose the SEE algorithm and derive explicit upper and lower bounds, decomposing $\mathbb{E}_{\nu}[\tau]$ into $\delta$-dependent and $\delta$-independent components. They show that when $m=1$, i.e. there is only one arm whose reward is greater than $\mu_{0}$, the gap between upper and lower bounds is at most polynomial in $\log K$ and $\log(1/\Delta_{0,1}^{2})$, implying near-optimality. For $m>1$, however, they comment the $\delta$-independent term $H(1)\log\!\bigl(K/\Delta_{0,1}^{2}\bigr)$ may be improvable, which is explicitly left as an open problem.